OpenStudy (anonymous):

Write Cos(3pi/2 - theta) in terms of cosine by using trig identities

jimthompson5910 (jim_thompson5910):

hint: cos(x-y) = cos(x)*cos(y) + sin(x)*sin(y)

jimthompson5910 (jim_thompson5910):

another hint: using the unit circle, you will see that cos(3pi/2) = 0 sin(3pi/2) = -1

OpenStudy (anonymous):

but that gets you -sin theta which isn't in terms of cosine

jimthompson5910 (jim_thompson5910):

well you could use the idea that cos(pi/2 - x) = sin(x), but that gets you back to the starting point in a way

jimthompson5910 (jim_thompson5910):

also, sin^2(x) + cos^2(x) = 1 so you could solve for sin(x)